Determine the current price (P0) of a constant growth stock that has a D1 of $2.15, a constant growth rate (g) of 5%, and a rate of return (rS) of 12%.

$40 $50 $30 $20

A company has an annual coupon bond issue that has a coupon rate of 3%, a current price of $918.89, and 10 years of maturity. Determine the bond’s YTM.

Cannot be determined without knowing its YTC. 4% 5% 3%

Other things being equal, which of the following types of bonds should have a lower yield to maturity?

All bonds should have the same yield to maturity.

A bond is a debenture.

A bond has no collateral.

A bond is secured by land.
Answer #1

1) $30

According to dividend discount model, current stock price is the present value of future dividends

Current stock price = Next year's dividend/(r-g)

r= required rate = 12%

g = Constant growth rate = 5%

Current stock price = 2.15/(0.12-0.05) = 30.714 ~ $30

2) 4%

Using a financial calculator

PV = 918.89

FV = 1000

PMT = 30 (3% coupon on face value $1000)

N = 10

cpt I/Y, we get I/Y= 4

Hence, YTM = 4.00%

3) A bond is secured by land.

Lower yield is associated with lower risk. A debenture or a bond with no collateral has higher risk and hence would have higher yield to maturity. A bond with land as collateral is relatively safer and hence has lower yield to maturity.